# AlabOS: A Python-based Reconfigurable Workflow Management Framework for Autonomous Laboratories

### v1

@@ 3.4 Data storage
Despite the logger’s ease of use, logged data can still become unstructured and scattered in the database. For example, each point of the real-time temperature of the furnace can be a single log document in the database. To make the data more accessible to researchers, we also provide a “result” field for each task. At the end of each task, the task can gather data generated during the run and return them as a Python dictionary. AlabOS stores all returned data in the result field. For example, one can return the temperature-time curve at the end of the heating task, which is especially useful when troubleshooting failed syntheses.
@@ 4 AlabOS in practice: orchestrating an autonomous laboratory for solid-state synthesis
Figure 5: The typical tasks and workflows and sample throughput in the A-Lab. (a) In a typical workflow, one powder sample can be heated either in a box furnace (Heating task) or in a tube furnace where the atmosphere is controlled (HeatingWithAtmosphere task). (b) A batched task graph with 16 samples, including eight samples sent to the box furnace and eight to the tube furnace. (c) The number of distinct samples processed by A-Lab by quarter between March 2023 and April 2024.
To set up the A-Lab workflows using AlabOS, a project folder is first created. A toml format configuration file is made to store all the connection information to the database, the notification service, and the message queue service. The devices’ driver codes are stored in a devices folder. Each type of device is defined in a Python object class inherited from BaseDevice to include some basic functions like connecting and status checking. Apart from these, we can define as many methods as needed to operate the device. We also define the associated sample position names in the device. They are read by AlabOS and recorded in the database when launching.
@@
Task name Capacity
Successfully
completed tasks
PowderDosing 16 248
Heating 8 233
ManualHeating 8 65
HeatingWithAtmosphere 4 25
RecoverPowder 1 1865
Diffraction 1 1845
Ending 1 1752
Table 1: The maximum capacity of each task and the number of completed tasks in A-Lab by the time of writing.
@@
By the time of writing, AlabOS has been driving A-Lab to synthesize and characterize over 2,500 distinct samples over a period of approximately one year, as shown in Fig .5c. The maximum number of samples submitted in one single day is 149 samples on Feb 9, 2024. The system is designed to be able to handle a large amount of sample submissions at a time. By dividing the experiments composed of many samples into modular tasks, AlabOS is able to schedule and manage experiments at a finer granularity, thus ensuring the high efficiency of completing the complicated workflows.
@@ 5 Conclusion and Outlook
We have outlined the development and application of AlabOS as an orchestration software for managing workflows in autonomous laboratories. This package is designed to be highly general and reconfigurable, which are both necessary features to meet the fast-changing requirements of a typical materials research program. With the independent task actor design and the resource occupation mechanism, human operators can define and submit workflows in AlabOS without concerning themselves with possible conflicts between concurrent tasks. The dashboard and notification system provide a general solution for human-machine interaction in the laboratory, which becomes especially useful when performing maintenance and error recovery jobs. With the public availability of AlabOS, we hope that researchers will no longer need to write complex workflow management codes but instead focus their time on developing high-level logic followed by autonomous laboratories for accelerated materials discovery and optimization.
@@ 6 Code availability
The code for the AlabOS is available at https://github.com/CederGroupHub/alabos. The version of the code employed for this study is v1.0.0.

### v2

@@ 3.4 Data storage
Despite the logger’s ease of use, logged data can still become unstructured and scattered in the database. For example, each point of the real-time temperature of the furnace can be a single log document in the database. To make the data more accessible to researchers, we also provide a “result” field for each task. At the end of each task, the task can gather data generated during the run and return them as a Python dictionary. AlabOS stores all returned data in the result field. For example, one can return the temperature-time curve at the end of the heating task, which is especially useful when troubleshooting failed syntheses. If a large file needs to be stored in the database, AlabOS also provides a “LargeResult” object to handle the large file with MongoDB GridFS backend. The large file will be stored as chunks, with a reference ID generated and linked to the result collection.
@@
3.5 Experiment Submission
In AlabOS, all the experiments are submitted through the dashboard server, where an API is exposed to receive all the experiment submissions. The format of the experiment is first validated to ensure the correct processing afterward. The validated experiment is then dumped into an experiment collection, where the experiment manager instance keeps polling and builds samples and task graphs according to the experiment specifications. The raw input format is in JSON format, which includes the sample and task information. Although it is easy for the system to parse the JSON format, it is usually not straightforward for the users to compose the input file directly. An experiment builder class is introduced to help users to define experiments in a Python script. When an experiment builder is created, the name and metadata of the experiment should be provided. With an “add_sample” method, the users can create and add samples to the experiments. Then the tasks are created and attached to the samples. The experiment builder will record the sequence of the tasks added to each sample and build the task graph with the dependency specified in the sequence. Finally, the experiment can be converted to a JSON format string and submitted to the submission API with a “submit” method. The users can also query the experiment status and results from the experiment information API.
The submission API and the experiment builder serve as a high-level abstraction to the autonomous laboratory. It resembles the queue management system in high-performance computing (HPC) clusters. When submitting jobs to HPC, the users do not need to pay attention to how and where their jobs are run. Similarly, when an experiment is submitted to AlabOS, the users can focus on designing experiments with proper parameters without knowing the details of how the experiment will be run.
The submission infrastructure of AlabOS opens up the possibility of conducting a close-loop experiment campaign inside an autonomous laboratory, where AI agents analyze the results and make plans for future experiments. With a structured input format that is sent to an API endpoint, there is no distinction between experiments submitted by humans and by AI. When designing a close-loop experiment campaign, the user can wrap the submission and result query script into an objective function with the inputs and outputs to be explored. The objective function can be directly used in various experiment planning software for self-driving laboratories like Chimera 59 and Altas 60.
@@ 4 AlabOS in practice: orchestrating an autonomous laboratory for solid-state synthesis
Figure 6: The typical tasks and workflows and sample throughput in the A-Lab. (a) In a typical workflow, one powder sample can be heated either in a box furnace (Heating task) or in a tube furnace where the atmosphere is controlled (HeatingWithAtmosphere task). (b) A batched task graph with 16 samples, including eight samples sent to the box furnace and eight to the tube furnace. (c) The number of distinct samples processed by A-Lab by quarter between March 2023 and July 2024.
To set up the A-Lab workflows using AlabOS, a project folder is first created. A toml format configuration file is made to store all the connection information to the database, the notification service, and the message queue service. The devices’ driver codes are stored in a devices folder. Each type of device is defined in a Python object class inherited from BaseDevice to include some basic functions like connecting and status checking. Apart from these, we can define as many methods as needed to operate the device. We also define the associated sample position names in the device. They are read by AlabOS and recorded in the database when launching. To build an automated lab for solid-state synthesis and powder XRD characterization, there are in total 16 types of devices with 28 device instances and 289 device-related sample positions. The full list of defined devices, their communication protocols, the number of associated sample positions, and numbers of handled exceptions are listed in Table .1. Most devices communicate with the control PC through Ethernet, with various protocols including HTTP, MODBUS, XML-RPC, etc. Each device (and control PC) is assigned to a unique IP within an intranet. The control PC can send commands to the devices by specifying the IP addresses of the devices. The box furnaces, as an exception, are connected directly to the control PC through its serial interfaces, with a COM port assigned to each furnace.
Device Type Physical Device Communication Protocol Number of Devices Number of Associated Sample Positions Per Device
BallDispenser Arduino HTTP 1 1
BoxFurnace Thermo Scientific F4805560 MODBUS over Serial 4 8
CapDispenser Arduino HTTP 1 3
CappingGripper Arduino HTTP 1 1
Diffractometer Malvern Panalytical Aeris Mineral Edition Socket 1 1
LabmanQuadrant Labman Powder Dosing System HTTP 4 48
ManualFurnace N/A N/A 4 8
RobotArmCharacterization Universal Robot 5e Socket & SSH 1 1
RobotArmFurnaces Universal Robot 5e Socket & SSH 1 1
Scale OHAUS Scout HTTP 1 1
Shaker Arduino HTTP 1 1
TransferRack N/A N/A 1 6
TubeFurnace MTI OTF1200X5ASD XML-RPC 4 4
VialDispenserRack N/A N/A 1 0
VialLabeler Reiner jetStamp 1025 Serial 1 1
XRDDispenserRack N/A N/A 1 0
Table 1: All the devices implemented in A-Lab within AlabOS framework. Each device type represents a Python class definition using AlabOS APIs. The physical devices used in A-Lab are listed. The communication protocol indicates how the device object communicates with the physical device in A-Lab. If the protocol is N/A, it indicates the device is not linked to any physical devices and only exists in AlabOS system to manage some status of the lab. For example, the XRDDispenserRack is used to monitor the number of clean XRD sample holders and will notify users when there are no more clean XRD holders. The number of devices and the associated sample positions for each device are also listed.
@@
Task name Capacity Number of tasks Error handling routines in definition % with exceptions % with unrecoverable exceptions
PowderDosing 16 333 1 1.20 1.20
Heating 8 403 2 4.96 4.22
ManualHeating 8 75 0 0 0
HeatingWithAtmosphere 4 27 0 22.22 7.41
RecoverPowder 1 2818 15 4.19 0.99
Diffraction 1 2658 11 4.74 1.20
Ending 1 2574 3 0.12 0.04
Table 2: The maximum capacity of each task and the number of tasks in A-Lab by the time of writing. The number of handled errors in each task, indicated by the try/except blocks with user inputs requested, is shown in the table. Among all the tasks, the ratios of tasks with exceptions and with unrecoverable exceptions are also shown in the last two columns.
@@
By the time of writing, AlabOS has been driving A-Lab to synthesize and characterize over 3,500 distinct samples over a period of approximately one and half years, as shown in Fig .6c. The maximum number of samples submitted in one single day was 149 samples on Feb 9, 2024. The system is designed to be able to handle a large amount of sample submissions at a time. By dividing the experiments composed of many samples into modular tasks, AlabOS is able to schedule and manage experiments at a finer granularity, thus ensuring the high efficiency of completing the complicated workflows.
To maintain smooth operation, A-Lab task definitions incorporate handling routines for common exceptions. The number of exceptions in each A-Lab task is listed in Table. 2. Depending on the complexity of the task, different exception-handling codes are inserted into task definitions. Among them, RecoverPowder and Diffraction require the most exception handling as many steps are involved in the operation. The number of exception-handling routines also aligns with the ratio of recovered exceptions in all the exceptions raised, as shown in Table. 2. RecoverPowder and Diffraction demonstrate a higher exception recovery rate of 76% and 75%, respectively. On the contrary, Heating and PowderDosing have a lower recovery rate of 15% and 0% due to limited exception handling. The lack of routines to address exceptions in these tasks can be attributed to the fact that some are due to hardware that is worn out, or to newly emerged software communication exceptions, which were not detected in the soak test. Over time, these tasks will suffer less from unrecoverable errors in the future with more exception-handling routines included.
@@ 5 Conclusion and Outlook
We have outlined the development and application of AlabOS as an orchestration software for managing workflows in autonomous laboratories. This package is designed to meet the fast-changing requirements of a typical materials research program. With the graph-based experiment input format, independent task actor design, and the resource occupation mechanism, human operators can define and submit workflows in AlabOS without concerning themselves with possible conflicts between concurrent tasks. The dashboard and notification system provide a general solution for human-machine interaction in the laboratory, which becomes especially useful when performing maintenance and error recovery jobs. With the public availability of AlabOS, we hope that researchers will no longer need to write complex workflow management codes but instead focus their time on developing high-level logic followed by autonomous laboratories for accelerated materials discovery and optimization. Only in this way, the autonomous laboratory can be more accessible to general researchers and become a more powerful tool for accelerating material discovery.
@@ 6 Code availability
The code for the AlabOS is available at https://github.com/CederGroupHub/alabos. The version of the AlabOS code employed for this study is v1.0.1. The online document is hosted at https://cedergrouphub.github.io/alabos/. The implementation of device communication driver is available at https://github.com/CederGroup/alab_control. The A-Lab-specific device and task definitions are available at https://github.com/CederGroupHub/alabos/tree/main/examples/alab_example.